### v2

# Differentially Private Sampling via Reveal-or-Obscure

Abstract
We introduce a differentially private (DP) algorithm called Reveal-or-Obscure (ROO) to generate a single representative sample from a dataset of $n$ i.i.d. observations from an unknown distribution. Unlike methods that add explicit noise to the estimated empirical distribution, ROO achieves $\epsilon$-differential privacy by choosing whether to “reveal" or “obscure" the empirical distribution with a fixed probability $q$. While our proposed mechanism is structurally identical to an algorithm proposed by Cheu and Nayak, we prove a strictly better bound on the sampling complexity than that established in their theorem. Building on this framework, we propose a novel generalized sampler called Data-Specific ROO (DS-ROO), where the obscuring probability $q$ is a function of the empirical distribution. We show that when the dataset contains enough samples from every element of the alphabet, DS-ROO can achieve $\epsilon$-DP while obscuring much less. In addition, we provide tight upper bounds on the utility of DS-ROO in terms of total variation distance. Our results show that under the same privacy budget, DS-ROO can achieve better utility than state-of-the-art private samplers and vanilla ROO, with total variation distance decaying exponentially in dataset size $n$.
I Introduction
The widespread use of sensitive data across various domains, including healthcare, finance, law enforcement, and social sciences, has heightened the importance of privacy-preserving data analysis. Consequently, there is a growing need for mechanisms that allow data analysis while minimizing individual privacy risks. One promising approach is the use of synthetic data that capture the statistical properties of the original data.
Differential Privacy (DP) [2, 3] has emerged as a sound framework for formalizing privacy guarantees across a range of applications, including data analysis. In essence, DP ensures that the output of an algorithm does not differ by much whether or not an individual’s data is included in the input.
The task of synthetic data generation is closely related to the broader problem of learning probability distributions, which has been studied under both central and local DP; for example, see [4, 5, 6, 7]. In the non-private setting, a learning algorithm approximates a distribution from which one can sample new data points that are representative of the original data. When privacy constraints are introduced, learning a distribution becomes significantly more challenging. In many practical cases, it may be sufficient to produce a small number of representative samples instead.
The task of privately releasing one sample—known as DP sampling—is easier than full-fledged learning, since it requires less information from the underlying distribution. Motivated by this, we introduce a novel DP sampling algorithm for discrete distributions on a finite alphabet. The key idea of our approach is to “obscure” the empirical distribution of the input dataset with a certain probability, or “reveal” it otherwise. Hence, we call our proposed algorithm reveal-or-obscure (ROO). ROO can be viewed as a form of randomized response, where it outputs a sample drawn from a mixture of two distributions. We further generalize this idea through a data-specific algorithm called Data-Specific ROO (DS-ROO), in which the probability of sampling from the uniform distribution is a function of the smallest empirical probability of the dataset. This adaptive design enables DS-ROO to obscure the empirical distribution less often when the dataset contains sufficiently many samples of each element, leading to significantly improved accuracy under the same privacy budget. Table I summarizes the total variation distance achieved by ROO, DS-ROO, and several existing private samplers, illustrating the utility gains obtained by our proposed mechanisms.
Main Contributions. We focus on the setting where we are given a dataset of $n$ samples drawn from an unknown distribution with support size $k$ and a privacy parameter $\epsilon$. In keeping with prior work, we define accuracy (utility) via the total variation distance. Our main contributions are as follows:
  • We propose ROO—a sampling algorithm that achieves differential privacy without explicitly perturbing the empirical distribution of the input dataset. We incorporate uncertainty in our algorithm by sampling from the uniform distribution with some fixed probability $q$, which we henceforth refer to as obscuring probability.
  • We prove that our proposed algorithm ROO achieves $\epsilon$-DP and $\alpha$-accuracy. Additionally, we show that ROO reduces the sampling complexity while achieving better privacy-utility trade-off than the state-of-the art [8, 9].
  • We also propose Data-Specific ROO (DS-ROO) as a technique to generalize ROO by making $q$, i.e., the probability of sampling from the uniform distribution, a function of the empirical distribution of the dataset. We prove that DS-ROO achieves $\epsilon$-DP and its utility is always at least as good as ROO’s. In other words, the data-specific obscuring probability is no larger than that of the vanilla ROO algorithm.
  • We provide theoretical upper bounds on the utility of DS-ROO, quantified by its total variation distance, under the assumption that each element of the alphabet appears in the true distribution with probability at least $\gamma$. In particular, we show that the total variation distance goes exponentially to zero as dataset size $n\to\infty$. Our bound on the utility of DS-ROO, asymptotic in $n$ as $n\to\infty$, is shown in Table I.
  • We demonstrate numerically that, for the same privacy guarantee, DS-ROO achieves significantly better utility than vanilla ROO as well as the state-of-the-art methods [8, 9].
TABLE I: Comparison of Accuracy of Private Sampling Mechanisms in Terms of Total Variation Distance with Parameters: $n$ (sample size), $k$ (alphabet size), $\epsilon$ (privacy budget), $\gamma$ (smallest probability of occurrence of an element in the true distribution). The $d_{TV}$ bounds hold for all values of $n$ except for the result for DS-ROO, which holds asymptotically as $n\to\infty$.
Raskhodnikova et al. [8] Cheu & Nayak [9] ROO (Ours) [1] DS-ROO (Ours)
$d_{\mathrm{TV}}$ $\dfrac{2k}{n\epsilon}$ $\dfrac{k-1}{n\epsilon+k-1}$ $\dfrac{1-\dfrac{1}{k}}{1+\dfrac{n}{k}(e^{\epsilon}-1)}$ $\dfrac{(k-1)e^{-2n\gamma^{2}+4\gamma m_{0}+o(1)}}{1+\dfrac{n}{k}(e^{\epsilon}-1)},\text{ where}$
$m_{0}=\left\lceil\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\}\right\rceil$
Related Work. The problem of differentially private sampling from unknown distributions is first investigated in [8]. Raskhodnikova et al. [8] provide the first known bounds with $(\epsilon,\delta)$-DP guarantees on the complexity of sampling from arbitrary distributions over a discrete alphabet. DP algorithms for sampling from higher dimensional distributions such as multivariate Gaussians are presented in [10]. Husain et al. [11] consider DP sampling in the local setting, where the central aggregator cannot be trusted and each user must produce a single data record privately. Private sampling has also been studied in the local model; recent works propose minimax-optimal mechanisms for all $f$-divergences [12], and leverage public data to improve utility under local DP [13]. A key focus of these efforts has been to reduce the sample complexity of DP-assured private sampling. A recent and concurrent work of Cheu and Nayak [9] independently proposes an algorithm structurally identical to ROO. However, our analysis establishes a strictly better sampling complexity bound in the same setting.
More generally, the problem of releasing a dataset in a differentially private manner has also been studied; for example, see [14, 15, 16, 17, 18]. Generating a single sample in a private manner is the first step towards releasing a larger synthetic dataset, and to this end, we focus on the former challenge in this paper.
The remainder of this paper is organized as follows. We review various theoretical concepts relevant to the problem of differentially private sampling in Section II. In Section III, we present our proposed private sampling algorithm ROO and prove that it achieves $\epsilon$-DP with $\alpha$-accuracy. In Section IV, we present our generalized sampler DS-ROO, prove that it achieves $\epsilon$-DP, and provide upper bounds on its utility. Where relevant, we provide proof sketches of our theorems. We present numerical comparison of the utility of various private sampling algorithms in Section V and show that DS-ROO achieves superior performance for the same privacy guarantee, and conclude in Section VI.
II Problem Setup
We begin by briefly reviewing some relevant definitions. We use uppercase letters, e.g., $X$, to denote random variables (RVs), and lowercase letters, e.g. $x$, for their instantiations. We assume that the dataset consists of $n$ RVs sampled from a finite alphabet of $k$ letters; without loss of generality, we take this alphabet to be $k]=\{1,2,\dots,k\}$. Let $\mathcal{P}$ be the class of probability distributions on $[k]$. Given a dataset $X^{n}=(X_{1},X_{2},\dots,X_{n})$ of $n$ i.i.d. observations from some unknown $P\in\mathcal{P}$, a randomized algorithm (privacy mechanism) $\mathcal{A}:\mathcal{X}^{n}\mapsto\mathcal{X}$ outputs a single sample from $\mathcal{X}$. Let $Y=\mathcal{A}(X^{n})$ be the random variable corresponding to the output of algorithm $\mathcal{A}$ given input $X^{n}$. The output $\mathcal{A}(X^{n})$ is drawn from a distribution $Q$ such that
$Q(y)=\sum_{x^{n}\in\mathcal{X}^{n}}\Pr\left\{\mathcal{A}(x^{n})=y|x^{n}\right\}\Pr\left\{X^{n}=x^{n}\right\}.$ (1)
The accuracy of $\mathcal{A}$ is measured by the closeness between $Q$ and $P$. We use the total variation distance, defined as
$d_{TV}(Q,P)=\frac{1}{2}\|Q-P\|_{1}=\frac{1}{2}\sum_{x\in\mathcal{X}}|Q(x)-P(x)|.$ (2)
We use the following definition of sampling accuracy, introduced in [19].
Definition 1 (Accuracy of Sampling).
An algorithm $\mathcal{A}$ is $\alpha$-accurate on a distribution $P$ if
$d_{TV}(Q,P)\leq\alpha.$ (3)
An algorithm is $\alpha$-accurate on a class $\mathcal{P}$ of distributions if it is $\alpha$-accurate on every $P\in\mathcal{P}$.
Two datasets $x^{n}$ and $\tilde{x}^{n}$ are considered neighbors, denoted $x^{n}\sim\tilde{x}^{n}$, if they differ by at most one entry. DP is defined with respect to all such neighboring datasets as follows.
Definition 2 (Differential Privacy [2]).
A randomized algorithm, or mechanism $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{Y}$ is considered $\epsilon$-differentially private ($\epsilon$-DP) if, for every pair of neighboring datasets $x^{n}\sim\tilde{x}^{n}\in\mathcal{X}^{n}$, and for $B\subseteq\mathcal{Y}$,
$\Pr\{\mathcal{A}(x^{n})\in B\}\leq e^{\epsilon}\Pr\{\mathcal{A}(\tilde{x}^{n})\in B\}.$ (4)
II-A State-of-the-art Algorithms
In [8], the authors present an achievable $\epsilon$-DP sampler which does the following:
  1. 1.
    Computes, for each $j\in[k]$, the empirical probability distribution $\hat{p}_{j}$,
  2. 2.
    Adds noise drawn from $\text{Lap}(\frac{2}{n\epsilon})$ to each count,
  3. 3.
    Uses an $L_{1}$ projection to restrict the noisy distribution to be a probability vector $\tilde{P}=(\tilde{p}_{1},\dots,\tilde{p}_{k})$, and
  4. 4.
    Outputs an element of $[k]$ sampled from $\tilde{P}$.
They show that this algorithm is $\alpha$-accurate with a sampling complexity
$n^{\prime}=\frac{2k}{\alpha\epsilon}.$ (5)
Cheu and Nayak [9] propose Subsampled Randomized Response (SubRR), which selects a data point uniformly at random from the input dataset and applies $k$-ary randomized response [20, 2]. In [9, Theorem 12], they show that their algorithm is $\alpha$-accurate with a reduced sampling complexity
$n^{\prime\prime}=\frac{1}{\alpha\epsilon}(k-1)(1-\alpha).$ (6)
In the following section, we prove that it is possible to achieve an $\epsilon$-DP and $\alpha$-accurate sampler with fewer samples than both $n^{\prime}$ and $n^{\prime\prime}$. Notably, for higher values of $\epsilon$, i.e., lower privacy, we gain an exponential reduction in the required number of samples. Moreover, [8] also establishes a lower bound on the sampling complexity as $\Omega\left(\frac{k}{\alpha\epsilon}\right)$ for a restricted range of $\epsilon\in(0,1]$.
III Reveal-or-Obscure (ROO)
Algorithm 1 presents our proposed private sampler ROO. We implement the idea of obscuring the empirical distribution $\hat{P}_{x^{n}}$ by sampling from the uniform distribution on $[k]$. However, we wish to do so with a small probability $q$, so that we do not deviate too much from the true distribution $P$. With probability 1q1-q, we simply choose a sample from the given dataset, i.e., we reveal $\hat{P}_{x^{n}}$. We refer to $q$ as the obscuring probability, since it quantifies how often ROO obscures the empirical distribution by outputting a uniformly distributed sample instead.
Algorithm 1 Reveal-or-Obscure (ROO)
1:Dataset $x^{n}=(x_{1},\dots,x_{n})$, alphabet size $k$, privacy budget $\epsilon$, obscuring probability $q$
2:Sample $y$
3:With probability $q$, choose $y\sim\mathrm{Unif}[1\!:\!k]$
4:Otherwise, sample $i\sim\mathrm{Unif}[1\!:\!n]$ and set $y\leftarrow x_{i}$
5:return $y$
The privacy and utility guarantees provided by Algorithm 1 are given by the following theorem.
Theorem 1.
Given $q$, ROO (Algorithm 1) is $\epsilon$-DP and $\alpha$-accurate for
$\epsilon$ $=\log\left(1+\frac{k(1-q)}{nq}\right),\text{ and}$ (7)
$\alpha$ $=q\left(1-\frac{1}{k}\right),$ (8)
from which we can solve for $q$ to obtain the sampling complexity as
$n=\frac{k(1-\alpha)-1}{\alpha(e^{\epsilon}-1)}.$ (9)
Remark 1.
For a given dataset size $n$, alphabet size $k$ and privacy budget $\epsilon$, the corresponding obscuring probability $q$ can be obtained by rearranging the privacy expression (7) in Theorem 1:
$q=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}.$ (10)
Proof:
The proof consists of two parts—a privacy analysis, which establishes the $\epsilon$-DP guarantee, and a utility analysis, which derives the $\alpha$-accuracy bound.
Privacy Analysis. Given $x^{n}$, Algorithm 1 chooses output $Y=y$ with probability
$\Pr\{Y=y|x^{n}\}$ $=\frac{q}{k}+(1-q)\hat{P}_{x^{n}}(y),$ (11)
where $\hat{P}_{x^{n}}(x)$ denotes the empirical probability of each $x\in[k]$,
$\hat{P}_{x^{n}}(x)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(x_{i}=x).$ (12)
For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, their corresponding empirical probabilities of one observation differ by at most $\frac{1}{n}$, i.e., for all $x\in\mathcal{X}$,
$\left|\hat{P}_{x^{n}}(x)-\hat{P}_{\tilde{x}^{n}}(x)\right|\leq\frac{1}{n}.$ (13)
By Definition 2, in order for ROO to satisfy $\epsilon$-DP, the following condition must hold for all possible $x^{n}\sim\tilde{x}^{n}$,
$\frac{\Pr\{Y=y|x^{n}\}}{\Pr\{Y=y|\tilde{x}^{n}\}}\leq e^{\epsilon}.$ (14)
We can write the left side of (14) as
$\frac{\Pr\{Y=y|x^{n}\}}{\Pr\{Y=y|\tilde{x}^{n}\}}$ $=\frac{\frac{q}{k}+(1-q)\hat{P}_{x^{n}}(y)}{\frac{q}{k}+(1-q)\hat{P}_{\tilde{x}^{n}}(y)}$ (15)
$\leq\sup_{p\in[0,1]}\frac{\frac{q}{k}+(1-q)\left(p+\frac{1}{n}\right)}{\frac{q}{k}+(1-q)p}$ (16)
$=\sup_{p\in[0,1]}1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}+(1-q)p}$ (17)
$\leq 1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}},$ (18)
where, in (15), we substitute (11); in (16), we use the property of the empirical probabilities stated in (13), denoting $\hat{P}_{\tilde{x}^{n}}(y)$ with $p$ for notational simplicity; in (18), the supremum is obtained when $p=0$.
Therefore, ROO satisfies $\epsilon$-DP guarantee if the right side of (18) is bounded by $e^{\epsilon}$, i.e.,
$1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}}\leq e^{\epsilon}.$ (19)
Rearranging (19), we obtain the privacy guarantee of Theorem 1 stated in (7).
Utility Analysis. For our proposed sampler ROO in Algorithm 1, the output distribution is
$Q(y)$ $=\sum_{x^{n}\in\mathcal{X}^{n}}\left(\frac{q}{k}+(1-q)\hat{P}_{x^{n}}(y)\right)\Pr\left\{X^{n}=x^{n}\right\}$ (20)
$=\frac{q}{k}\sum_{x^{n}\in\mathcal{X}^{n}}\Pr\left\{X^{n}=x^{n}\right\}$
$\quad\quad+(1-q)\sum_{x^{n}\in\mathcal{X}^{n}}\hat{P}_{x^{n}}(y)\Pr\left\{X^{n}=x^{n}\right\}$ (21)
$=\frac{q}{k}+(1-q)\mathbb{E}_{X^{n}}[\hat{P}_{X^{n}}(y)]$ (22)
$=\frac{q}{k}+(1-q)P(y),$ (23)
where (23) follows from the fact that $\hat{P}_{x^{n}}$ is the empirical distribution of a dataset sampled from $P$. The total variation distance between the discrete distributions $Q$ and $P$ is thus
$d_{TV}(Q,P)$ $=\frac{1}{2}\sum_{y\in\mathcal{X}}|Q(y)-P(y)|$ (24)
$=\frac{q}{2}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-P(y)\right|$ (25)
$=q\times d_{TV}\left(\frac{1}{k},P(y)\right)$ (26)
$\leq q\times\max_{P(y)}d_{TV}\left(\frac{1}{k},P(y)\right),$ (27)
where we rearrange and substitute (23) to obtain the expression in (25). Note that, for a convex objective function, the maximum is achieved at its corner points. Hence, the distribution $P$ that maximizes the $d_{TV}$ between the uniform distribution on $[k]$ and the input distribution $P$ must be one of the corner points of the $k$-dimensional probability simplex $\mathcal{P}$, e.g. $P=\{1,0,\dots,0\}$. The maximum $d_{TV}$ is then computed as
$\max_{P(y)}d_{TV}\left(\frac{1}{k},P(y)\right)$ $=\max_{P(y)}\frac{1}{2}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-P(y)\right|=1-\frac{1}{k}.$ (28)
Substituting this maximum objective value into (27), we have
$d_{TV}(Q,P)$ $\leq q\left(1-\frac{1}{k}\right).$ (29)
Comparing this result with Definition 1, we obtain the utility of Theorem 1, stated in (8). We can then express $q$, i.e., the probability of sampling from the uniform distribution, in terms of $\alpha$ and $k$,
$q=\frac{\alpha}{1-\frac{1}{k}}=\frac{k\alpha}{k-1}.$ (30)
Substituting the above into (19) and rearranging, we obtain sample complexity in (9). ∎
The expression in (9) reveals that Algorithm 1 requires substantially fewer samples than previous mechanisms. The following lemma formalizes this comparison.
Lemma 1.
For any $k\geq 2,\epsilon>0$, and $\alpha\in\left(0,1-\frac{1}{k}\right)$, the sampling complexity of Algorithm 1 is lower than that of [8] and [9].
The proof of Lemma 1 is given in Appendix A. In fact, the sampling complexity of Algorithm 1 is exponentially better in terms of $\epsilon$ compared to that of Raskhodnikova et al. [8] that we summarize in (5), and Cheu and Nayak [9] summarized in (6). This would appear to violate the lower bound of [8] on the sampling complexity; the reason it does not is that their lower bound only applies for $\epsilon<1$.
IV Data-Specific Reveal-or-Obscure (DS-ROO)
For Algorithm 1 in Section III, we fix $q$ in order to achieve $\epsilon$-DP for any possible dataset. From (18) in the proof of Theorem 1, we observe that the supremum of ratio of probabilities is achieved by setting $p=0$, where $p$ represents the empirical probability of one of the elements of the alphabet. This corresponds to the case where one of two neighboring datasets under consideration is entirely missing an element of the alphabet, but this element is present in the other dataset. In this section, we show that for datasets where each element of the alphabet appears reasonably often, it is possible to sample more efficiently than Algorithm 1. By making the obscuring probability $q$ a function of the smallest empirical probability of the dataset, the accuracy of the sampler can be improved for the same privacy. Let $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor\right\}$ denote the smallest number of times an element of $[k]$ appears in dataset $x^{n}$. That is, $m=n\cdot\min_{x\in[k]}\hat{P}_{x^{n}}(x)$. We note that since $X^{n}$ is a random vector, more generally, one can write a random variable $M$ that is the same function applied to $X^{n}$, for which $m$ is an instantiation.
The modified private sampler—which we call the data-specific reveal-or-obscure (DS-ROO)—is presented in Algorithm 2 and requires a one-time computation of $q_{m}$ which is outlined in steps 1–14. In DS-ROO, the data-specific obscuring probability $q_{m}$ is determined from the value of $m$ associated with the given dataset $x^{n}$. Algorithm 2 first computes the obscuring probability value $q_{m}$ via a recursive procedure; at first sight, it may not be obvious why this procedure is the way it is—its correctness will be demonstrated in the proof that follows. In steps 15–18, Algorithm 2 computes the minimum empirical probability $m$ from the input dataset $x^{n}$, then samples using the pre-computed $q_{m}$ as the obscuring probability. [^1] In fact, the computation of $(q_{m}:m\in\{0,\ldots,\lfloor\frac{n}{k}\rfloor\})$ can be done offline and only once given $k,n,\epsilon$.
Algorithm 2 Data-specific ROO (DS-ROO)
1:Dataset $x^{n}=(x_{1},\dots,x_{n})$, alphabet size $k$, privacy budget $\epsilon$
2:Sample $y$
3:Set $u^{\prime}\leftarrow-1+\frac{1}{k}-\frac{1}{n}$
4:Set $v^{\prime}\leftarrow e^{\epsilon}\!\left(\frac{1}{k}-1\right)$
5:Set $w^{\prime}\leftarrow e^{\epsilon}-1-\frac{1}{n}$
6:Compute $q_{0}\leftarrow\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}$
7:for $j=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor$ do
8:  $u_{j}\leftarrow-\frac{j}{n}+\frac{1}{k}-\frac{1}{n}$
9:  $v_{j}\leftarrow e^{\epsilon}\!\left(\frac{1}{k}-\frac{j}{n}\right)$
10:  $w_{j}\leftarrow-\frac{1}{n}-\frac{j}{n}+\frac{j}{n}e^{\epsilon}$
11:  if $j<\frac{n}{k}$ then
12:   $q_{j}\leftarrow\max\!\left\{\,0,\;\frac{u_{j}}{v_{j}}q_{j-1}-\frac{w_{j}}{v_{j}},\;\frac{v^{\prime}}{u^{\prime}}q_{j-1}+\frac{w^{\prime}}{u^{\prime}}\,\right\}$
13:  else if $j=\frac{n}{k}$ then
14:   $q_{j}\leftarrow\max\!\left\{\,0,\;\frac{v^{\prime}}{u^{\prime}}q_{j-1}+\frac{w^{\prime}}{u^{\prime}}\,\right\}$
15:  end if
16:end for
17:Compute $m\leftarrow n\cdot\displaystyle\min_{x}\hat{P}_{x^{n}}(x)$
18:With probability $q_{m}$, choose $y\sim\mathrm{Unif}[1\!:\!k]$
19:Otherwise, sample $i\sim\mathrm{Unif}[1\!:\!n]$ and set $y\leftarrow x_{i}$
20:return $y$
Figure 1: Plot of $q_{m}$ as a function of $m$ for fixed $k$ and $n$, showing changes under different privacy budgets $\epsilon$.
Notably, when $m=0$, the corresponding $q_{0}$ is equivalent to that of Algorithm 1. That is, the case of a dataset that is missing an element of the alphabet ($m=0$) is the worst-case scenario, and so in this case DS-ROO behaves identically to ROO. However, we will show both analytically and empirically that in other cases, DS-ROO can do much better than vanilla ROO.
Fig. 1 shows the plot of $q_{m}$ vs $m$ for a fixed $k=10,n=1000$, and several different values of the privacy parameter $\epsilon$. We observe that $q_{m}$ is non-increasing in $m$. As $m$ increases, the empirical distribution of the dataset gets closer to the uniform distribution, and $q_{m}$ approaches zero. Thus, DS-ROO is less likely to obscure the empirical distribution for larger $m$. In high privacy regimes, $q_{m}$ tends to decrease much slower. On the other hand, for larger $\epsilon$, i.e., low privacy, $q_{m}$ goes to zero much faster. The privacy guarantee provided by DS-ROO is given by the following theorem.
Theorem 2.
DS-ROO (Algorithm 2) is $\epsilon$-differentially private.
The complete proof of Theorem 2 is given in Appendix B. We provide a proof sketch below.
Proof:
For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, let $m$ and $\tilde{m}$ denote the smallest number of times any element of $[k]$ appears in $x^{n}$ and $\tilde{x}^{n}$, respectively. Since the two datasets differ in exactly one entry, there are three possible relationships between $m$ and $\tilde{m}$: (1) $\tilde{m}=m$, (2) $\tilde{m}=m+1$, or (3) $\tilde{m}=m-1$. Considering each of these cases separately and applying Definition 2, we derive the conditions under which DS-ROO satisfies $\epsilon$-DP as
$(u_{m}-v_{m})q_{m}$ $\geq w_{m},$ $\text{for all }m,$ (31)
$u_{m}q_{m+1}$ $\leq v_{m}q_{m}+w_{m},$ $\text{for }m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (32)
$u_{m}q_{m-1}$ $\leq v_{m}q_{m}+w_{m},$ $\text{for }m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor,$ (33)
$(u^{\prime}-v^{\prime})q_{m}$ $\leq w^{\prime},$ $\text{for all }m,$ (34)
$u^{\prime}q_{m+1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $\text{for }m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (35)
$u^{\prime}q_{m-1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $\text{for }m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (36)
where $u^{\prime},v^{\prime},w^{\prime},u_{m},v_{m}$, and $w_{m}$ are as defined in Algorithm 2. Additionally, for $q_{m}$ to be a valid probability, we must also have $0\leq q_{m}\leq 1$ for all $m$. We show that the following expression for $q_{m}$ satisfies the above conditions:
$q_{m}=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}},\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}\right\}.$ (37)
Here, the first term ensures $q_{m}$ is non-negative. The two remaining terms are obtained by rearranging (33) and (35). The quantity $q_{0}$ corresponds to the obscuring probability of the worst-case setting ($m=0$) and is equal to the $q$ value used in the original version of ROO, given in (10).
We subsequently show that:
  1. 1.
    $q_{m}$ is non-increasing in $m$.
  2. 2.
    The function $q_{m}$ in (37) satisfies the conditions in (31), (32), (34), and (36).
Therefore, Algorithm 2 is $\epsilon$-differentially private. ∎
The accuracy or utility guarantee of Algorithm 2 is given by the following theorem.
Theorem 3.
Assume the distribution $P$ satisfies $P(x)\geq\gamma$ for all $x\in[k]$, where $\gamma>0$. Given $k\geq 2$, $\epsilon>0$, and
$q_{0}=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)},$ (38)
DS-ROO (Algorithm 2) satisfies
$d_{TV}(P,Q)\leq\bigl(q_{m_{0}}+q_{0}\delta\bigr)\!\left(1-\frac{1}{k}\right),$ (39)
where $\delta=ke^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}$, and $q_{m_{0}}$ is the obscuring probability produced by Algorithm 2 at $m_{0}$. Furthermore, if there exists an integer $m_{0}$ such that $n,k,\epsilon,\gamma,m_{0}$ satisfy
$m_{0}$ $\geq\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\},$ (40)
$n$ $\geq\max\left\{\frac{m_{0}}{\gamma},\frac{k-1+\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}}{2(e^{\epsilon}-1)^{2}},2k\right\},$ (41)
then DS-ROO (Algorithm 2) is $\alpha$-accurate with
$\alpha=q_{0}(k-1)e^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}.$ (42)
The complete proof of Theorem 3 is given in Appendix C. We provide a proof sketch below.
Proof:
Lemma 8 shows that for any $\gamma^{\prime}<\gamma$,
$\Pr\!\left\{M<n\gamma^{\prime}\right\}\leq\delta=ke^{-2n(\gamma-\gamma^{\prime})^{2}},$ (43)
where $M=n\cdot\min_{x\in[k]}\hat{P}_{X^{n}}(x)$ is a random variable denoting the smallest empirical count in the dataset. Thus, the event $\{M<m_{0}\}$ with $m_{0}=n\gamma^{\prime}$ becomes exponentially unlikely as $n$ grows.
The utility of DS-ROO is determined by the total variation distance between the output distribution $Q$ of Algorithm 2 and the true distribution $P$. We can decompose $Q$ according to whether $M\geq m_{0}$ or $M<m_{0}$. Conditioned on $M\geq m_{0}$, the non-increasing obscuring probability $q_{M}$ is at most $q_{m_{0}}$, which implies
$\sum_{y}|Q(y)-P(y)|\leq 2q_{m_{0}}\left(1-\frac{1}{k}\right).$ (44)
When $M<m_{0}$, the mechanism behaves no worse than the ROO sampler, which implies
$\sum_{y}|Q(y)-P(y)|\leq 2q_{0}\left(1-\frac{1}{k}\right)\delta.$ (45)
Combining the two cases, we obtain the total variation distance
$d_{TV}(P,Q)\leq\bigl(q_{m_{0}}+q_{0}\delta\bigr)\!\left(1-\frac{1}{k}\right).$ (46)
The upper bound in (46) highlights the improvement on the utility of DS-ROO over the vanilla ROO algorithm. For ROO, the total variation distance is at most $q\left(1-\frac{1}{k}\right)$, where the obscuring probability $q$ is fixed and independent of the dataset. In contrast, the upper bound in (46) is smaller since DS-ROO operates with a reduced obscuring probability $q_{m}\leq q_{m_{0}}$, and the contribution of $q_{0}$ is attenuated by $\delta$. Note that, $q_{m_{0}}=0$ is a sufficient condition to eliminate the first term in (46). We show in Lemma 9 that $q_{m_{0}}=0$ whenever $m_{0}$ and $n$ satisfy the conditions in (40) and (41). Then the first term in (46) vanishes, and using $\delta=ke^{-2n(\gamma-\gamma^{\prime})^{2}}$ and $\gamma^{\prime}=\frac{m_{0}}{n}$ we obtain
$d_{TV}(P,Q)\leq\alpha=q_{0}(k-1)e^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}.$ (47)
Corollary 1.
For any distribution $P$ satisfying $P(x)\geq\gamma$ for all $x\in[k]$, and any $m_{0}$ satisfying (40), as $n\to\infty$, DS-ROO (Algorithm 2) achieves $\epsilon$-DP and $\alpha$-accuracy with
$\alpha$ $=q_{0}(k-1)e^{-2n\gamma^{2}+4\gamma m_{0}+o(1)}.$ (48)
This implies that the total variation distance goes exponentially to zero as $n\to\infty$.
V Numerical Results
In this section, we empirically compare the utility of DS-ROO with vanilla ROO and the state-of-the-art private sampling algorithms under same parameters. Recall from Definition 1 that the utility, i.e., accuracy, is measured in terms of total variation distance, and it is upper bounded by $\alpha$. Hence, smaller values of $\alpha$ corresponds to more accurate sampling. For the samplers of [8] and [9], we use the corresponding upper bounds on $d_{TV}$ as stated in Table I. From Theorem 1, we have the utility of ROO
$d_{TV}^{\text{ROO}}=\frac{1-\frac{1}{k}}{1+\frac{n}{k}(e^{\epsilon}-1)}.$ (49)
To evaluate the utility of DS-ROO for given values of $\gamma$, $n$, $\epsilon$, and $k$, we use the upper bound in (46) on the total variation distance. We compute $q_{m}$ for each $m$, and then evaluate (46) for each $m_{0}\in[1,n\gamma)$, which define as $d_{TV}(m_{0})$. Since DS-ROO reduces to vanilla ROO where all the assumptions made for the data-specific regime do not hold, we compute the best upper bound on the total variation of DS-ROO as:
$d_{TV}^{\text{DS-ROO}}(n)=\min\left\{d_{TV}^{\text{ROO}},\min_{m_{0}\in[1,n\gamma)}d_{TV}(m_{0})\right\}.$ (50)
Figure 2 shows the comparison of utility across a range of parameter settings. The top row shows results for fixed $k$ and $\epsilon$ while varying $\gamma$, and the bottom row shows results for fixed $\epsilon$ and $\gamma$ while varying $k$. Across all settings, ROO and DS-ROO outperform prior methods, with DS-ROO consistently achieving equal or better utility than vanilla ROO. For sufficiently large $\gamma$, the total variation distance for DS-ROO decays exponentially in $n$. We also observe that this exponential improvement is robust to variations in $\epsilon$ and $k$.
Figure 2: Comparison of accuracy in terms of $d_{TV}$ vs dataset size $n$ for various differentially private sampling algorithms. Top row: fixed $k,\epsilon$, varying $\gamma$; bottom row: fixed $\epsilon,\gamma$, varying $k$.
VI Conclusion
In this work, we study differentially private sampling algorithms for discrete distributions on a finite alphabet. For the algorithm we call Reveal-or-Obscure (ROO), we establish a strictly lower sampling complexity bound than was known in the literature. In addition, we propose a novel generalization—Data-Specific Reveal-or-Obscure (DS-ROO)—which adaptively selects the probability of obscuring the empirical distribution of the dataset to improve the privacy-utility trade-off. We prove that DS-ROO preserves the same $\epsilon$-DP guarantee as ROO while achieving better accuracy. Our theoretical bounds are supported by empirical evaluation demonstrating that DS-ROO outperforms existing private samplers when the underlying distribution is not in the worst-case configuration. For future work, we aim to explore the practicality of our proposed methods for more complex distribution classes, as well as for generating multiple samples.
Appendix A Proof of Lemma 1
We first compare the sampling complexity of ROO with that of [8]. From (9) in Theorem 1, we have the sampling complexity of ROO,
$n=\frac{k(1-\alpha)-1}{\alpha(e^{\epsilon}-1)}=\frac{k\left(1-\alpha-\frac{1}{k}\right)}{\alpha(e^{\epsilon}-1)}$ (51)
$\quad=\frac{2k}{\alpha\epsilon}\frac{\epsilon}{2(e^{\epsilon}-1)}\left(1-\frac{1}{k}-\alpha\right)<\frac{2k}{\alpha\epsilon}=n^{\prime}.$ (52)
As $\epsilon$ increases, the term $e^{\epsilon}-1$ in the denominator of (9) grows exponentially, resulting in significantly lower $n$ compared to $n^{\prime}$. Secondly, we compare the sampling complexity of ROO with that of [9]. Subtracting  (6) from (9), we have
$\frac{k(1-\alpha)-1}{\alpha(e^{\epsilon}-1)}-\frac{(k-1)(1-\alpha)}{\alpha\epsilon}$
$\quad\quad=\frac{k(1-\alpha)-1+\alpha-\alpha}{\alpha(e^{\epsilon}-1)}-\frac{(k-1)(1-\alpha)}{\alpha\epsilon}$ (53)
$\quad\quad=\frac{(k-1)(1-\alpha)-\alpha}{\alpha(e^{\epsilon}-1)}-\frac{(k-1)(1-\alpha)}{\alpha\epsilon}$ (54)
$\quad\quad=\frac{\epsilon(k-1)(1-\alpha)-\alpha\epsilon-(e^{\epsilon}-1)(k-1)(1-\alpha)}{\alpha\epsilon(e^{\epsilon}-1)}$ (55)
$\quad\quad=\frac{(\epsilon-e^{\epsilon}+1)(k-1)(1-\alpha)-\alpha\epsilon}{\alpha\epsilon(e^{\epsilon}-1)}<0,$ (56)
where (56) follows from the fact that $1+\epsilon<e^{\epsilon}$ for $\epsilon>0$. Therefore, despite the similar structure of ROO and SubRR, ROO requires fewer samples to achieve $\epsilon$-DP.
Appendix B Proof of Theorem 2
Since $m$ denotes the smallest number of times an element in the alphabet $[k]$ appears in a dataset $x^{n}$, it is at most $\left\lfloor\frac{n}{k}\right\rfloor$ when the empirical distribution is uniform. For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, there are three possible values of $\tilde{m}$: $(1)$ $\tilde{m}=m$, when the different entry in $\tilde{x}^{n}$ is a different element in $[k]$, hence it does not affect the minimum count, $(2)$ $\tilde{m}=m+1$, and $(3)$ $\tilde{m}=m-1$. In order to understand how the behavior of $q_{m}$ changes with $m$, we first analyze the likelihood ratio of the output distributions. Recall from Definition 2 that, for Algorithm 2 to achieve $\epsilon$-DP, the following condition must hold for all possible $x^{n}\sim\tilde{x}^{n}$ pairs:
$\max_{x}\frac{\frac{q_{\tilde{m}}}{k}+(1-q_{\tilde{m}})\hat{P}_{\tilde{x}^{n}}(x)}{\frac{q_{m}}{k}+(1-q_{m})\hat{P}_{x^{n}}(x)}\leq e^{\epsilon}.$ (57)
Adding and subtracting $\hat{P}_{x^{n}}(x)$ from $\hat{P}_{\tilde{x}^{n}}(x)$ in the numerator, the condition becomes
$\max_{x}\frac{\frac{q_{\tilde{m}}}{k}+(1-q_{\tilde{m}})(\hat{P}_{\tilde{x}^{n}}(x)-\hat{P}_{x^{n}}(x)+\hat{P}_{x^{n}}(x))}{\frac{q_{m}}{k}+(1-q_{m})\hat{P}_{x^{n}}(x)}\leq e^{\epsilon}.$ (58)
The ratio is maximized when the numerator is maximized. Using (13), we can rewrite the condition as
$\frac{\frac{q_{\tilde{m}}}{k}+(1-q_{\tilde{m}})(\frac{1}{n}+\hat{P}_{x^{n}}(x))}{\frac{q_{m}}{k}+(1-q_{m})\hat{P}_{x^{n}}(x)}\leq e^{\epsilon}$ (59)
$\Rightarrow\hat{P}_{x^{n}}(x)\left(1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})\right)$
$\quad\leq e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-\left(1-q_{\tilde{m}}\right)\frac{1}{n}.$ (60)
If (60) is satisfied at the two endpoints of $\hat{P}_{x^{n}}$, then the condition holds for all $\hat{P}_{x^{n}}$. Now, substituting $\min_{x}\hat{P}_{x^{n}}(x)=\frac{m}{n}$ into (60), we require
$\frac{m}{n}\left(1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})\right)\leq e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-\left(1-q_{\tilde{m}}\right)\frac{1}{n}.$ (61)
Let
$u^{\prime}$ $=-1+\frac{1}{k}-\frac{1}{n},$ (62)
$v^{\prime}$ $=e^{\epsilon}\!\left(\frac{1}{k}-1\right),$ (63)
$w^{\prime}$ $=e^{\epsilon}-1-\frac{1}{n},$ (64)
$u_{m}$ $=-\frac{m}{n}+\frac{1}{k}-\frac{1}{n},$ (65)
$v_{m}$ $=e^{\epsilon}\!\left(\frac{1}{k}-\frac{m}{n}\right),$ (66)
$w_{m}$ $=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}.$ (67)
Considering the three cases of $\tilde{m}$ and simplifying (61), we have the following conditions:
$(u_{m}-v_{m})q_{m}$ $\leq w_{m},$ $\text{for all }m,$ (68)
$u_{m}q_{m+1}$ $\leq v_{m}q_{m}+w_{m},$ $m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (69)
$u_{m}q_{m-1}$ $\leq v_{m}q_{m}+w_{m},$ $m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (70)
Similarly, substituting $\max_{x}\hat{P}_{x^{n}}(x)=1$ into (60), we require
$1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})\leq e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-\left(1-q_{\tilde{m}}\right)\frac{1}{n}.$ (71)
Considering the three cases of $\tilde{m}$ and simplifying (71), we have three more conditions:
$(u^{\prime}-v^{\prime})q_{m}$ $\leq w^{\prime},$ $\text{for all }m,$ (72)
$u^{\prime}q_{m+1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (73)
$u^{\prime}q_{m-1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (74)
Therefore, for Algorithm 2 to achieve $\epsilon$-DP, the function $q_{m}$ must satisfy (68)–(70) and (72)–(74). Now, recall the recursive definition of $q_{m}$ from Algorithm 2:
$q_{m}$ $=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}},\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}\right\}$ (75)
$=\max\left\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\right\}$ (76)
where we denote $f_{1}(m,q_{m-1})=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}$ and $f_{2}(q_{m-1})=\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}$. By inspection, (75) ensures that (70) and (73) are satisfied. Note that, since $q_{m}$ is a probability, we must have $0\leq q_{m}\leq 1$ for all $m$. The first term in (75) ensures that $q_{m}\geq 0$. The initial value $q_{0}$ is defined as in (10) is at most $1$; as we will show, $q_{m}$ is non-increasing in $m$. This ensures that $q_{m}\leq 1$.
The remainder of the proof is structured as follows:
  • Lemma 2 shows that (75) satisfies the condition in (68).
  • Lemma 3 shows that $f_{1}(m,q_{m-1})\leq q_{m-1}$.
  • Lemma 4 shows that $f_{2}(q_{m-1})\leq q_{m-1}$.
  • By Lemmas 3 and 4, we have
    $q_{m}=\max\left\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\right\}\leq\max\{0,q_{m-1}\}=q_{m-1}$ (77)
    where the last equality follows from the fact that $q_{m-1}\geq 0$. This proves that $q_{m}$ is non-increasing in $m$. This is used in the subsequent lemmas.
  • Lemma 5 shows that (75) satisfies the condition in (69).
  • Lemma 6 shows that (75) satisfies the condition in (72).
  • Lemma 7 shows that (75) satisfies the condition in (74).
B-A Auxiliary Lemmas
Lemma 2.
For $\epsilon>0,k\geq 2,n>k$, $m\in\left\{0,1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor\right\}$, and $u_{m},v_{m}$ and $w_{m}$ as defined in (65)–(67), the function $q_{m}$ satisfies the condition
$q_{m}\geq\frac{w_{m}}{u_{m}-v_{m}}.$ (78)
Proof:
Recall the inequality condition in (68),
$(u_{m}-v_{m})q_{m}\leq w_{m},$ (79)
for all $m$. Since $u_{m}-v_{m}<0$ for all $m$, the inequality in 79 is equivalent to
$q_{m}\geq\frac{w_{m}}{u_{m}-v_{m}}.$ (80)
Now, this condition is valid when the right side is positive. Since the denominator is always negative, this happens only when the numerator is also negative, i.e.,
$w_{m}=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}<0,$ (81)
which happens whenever
$m<\frac{1}{e^{\epsilon}-1}.$ (82)
Let
$t_{m}=\frac{w_{m}}{u_{m}-v_{m}}.$ (83)
Dividing the numerator and denominator by $v_{m}$ and rearranging, we have
$\left(\frac{u_{m}}{v_{m}}-1\right)t_{m}=\frac{w_{m}}{v_{m}}.$ (84)
We observe that
$t_{0}=\frac{w_{0}}{u_{0}-v_{0}}=\frac{-\frac{1}{n}}{\frac{1}{k}-\frac{1}{n}-\frac{e^{\epsilon}}{k}}=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}=q_{0}.$ (85)
Moreover, substituting the expressions for $u_{m},v_{m}$ and $w_{m}$ into (83), we have
$t_{m}$ $=\frac{\frac{m}{n}(e^{\epsilon}-1)-\frac{1}{n}}{\left(\frac{m}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (86)
$=\frac{\frac{m}{n}(e^{\epsilon}-1)-\frac{1}{n}-\frac{1}{k}(e^{\epsilon}-1)+\frac{1}{k}(e^{\epsilon}-1)}{\left(\frac{m}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (87)
$=\frac{\left(\frac{m}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}+\frac{1}{k}(e^{\epsilon}-1)}{\left(\frac{m}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (88)
$=1+\frac{\frac{1}{k}(e^{\epsilon}-1)}{\left(\frac{m}{n}-\frac{1}{k}\right)(e^{\epsilon}-1)-\frac{1}{n}}$ (89)
$=1-\frac{\frac{1}{k}(e^{\epsilon}-1)}{\left(\frac{1}{k}-\frac{m}{n}\right)(e^{\epsilon}-1)+\frac{1}{n}}.$ (90)
As $m$ increases, (90) becomes smaller due to a larger fraction being subtracted from $1$. Thus, $t_{m}$ is decreasing in $m$. Now, we prove that $q_{m}\geq t_{m}$ for the base case and apply induction.
Base case: For $m=1$, we have
$q_{1}$ $=\frac{u_{1}}{v_{1}}q_{0}-\frac{w_{1}}{v_{1}}$ (91)
$=\frac{u_{1}}{v_{1}}t_{0}-\left(\frac{u_{1}}{v_{1}}-1\right)t_{1}$ (92)
$=\frac{u_{1}}{v_{1}}(t_{0}-t_{1})+t_{1}$ (93)
$>t_{1},$ (94)
since $\frac{u_{1}}{v_{1}}>0$, and $t_{m}$ is strictly decreasing in $m$.
Induction step: Assume $q_{m}>t_{m}$ for any $m$. Then we have
$q_{m+1}$ $=\frac{u_{m+1}}{v_{m+1}}q_{m}-\frac{w_{m+1}}{v_{m+1}}$ (95)
$=\frac{u_{m+1}}{v_{m+1}}q_{m}-\left(\frac{u_{m+1}}{v_{m+1}}\right)t_{m+1}$ (96)
$>\frac{u_{m+1}}{v_{m+1}}t_{m}-\left(\frac{u_{m+1}}{v_{m+1}}\right)t_{m+1}$ (97)
Rearranging the right side of (97), we have
$q_{m+1}$ $>\frac{u_{m+1}}{v_{m+1}}(t_{m}-t_{m+1})+t_{m+1}.$ (98)
Since $t_{m}$ is strictly decreasing in $m$ and the coefficient of $t_{m}-t_{m+1}$ is positive, $q_{m+1}>t_{m+1}$. Thus, for $m\in\left\{0,1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor\right\}$, $q_{m}$ satisfies (78). ∎
Lemma 3.
For $\epsilon>0,k\geq 2,n>k$, and $m<\frac{n}{k}$
$f_{1}(m,q_{m-1})\leq q_{m-1}.$ (99)
Proof:
We may write
$f_{1}(m,q_{m-1})-q_{m-1}$ $=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}-q_{m-1}$ (100)
$=\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}.$ (101)
Thus, to prove that $f_{1}(m,q_{m-1})\leq q_{m-1}$, it suffices to show that the right-hand side of (101) is non-positive. Substituting (65) and (66) into the coefficient of $q_{m-1}$ in (101), we have
$\frac{u_{m}}{v_{m}}-1$ $=\frac{-\frac{m}{n}+\frac{1}{k}-\frac{1}{n}}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}-1$ (102)
$=\frac{\left(\frac{1}{k}-\frac{m}{n}\right)\frac{1}{n}-e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}$ (103)
$=\frac{-\frac{1}{n}-\left(\frac{1}{k}-\frac{m}{n}\right)(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}$ (104)
$<0,$ (105)
for all $m<\frac{n}{k}$. Similarly, substituting (67) and (66) into the second term in (101), we have
$\frac{w_{m}}{v_{m}}$ $=\frac{-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}$ (106)
$=\frac{-\frac{1}{n}+\frac{m}{n}(e^{\epsilon}-1)}{e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)}$ (107)
$>0,$ (108)
for $\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor<m<\frac{n}{k}$. Since $q_{m}$ has a lower bound 0 in this range, the right side of (101) is
$\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}<0.$ (109)
It still remains to show that the right side is negative for $m<\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$. We observe from Lemma 2 that in this range of $m$, $q_{m}$ has a non-zero and positive lower bound $t_{m}$ given by the right-hand side of (78). Substituting this bound to the right side of (101), we have
$\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}$ $\leq\left(\frac{u_{m}}{v_{m}}-1\right)t_{m-1}-\frac{w_{m}}{v_{m}}.$ (110)
Since $\frac{u_{m}}{v_{m}}-1<0$ and $t_{m}$ is strictly decreasing in $m$, we can write
$\left(\frac{u_{m}}{v_{m}}-1\right)t_{m-1}-\frac{w_{m}}{v_{m}}$ $\leq\left(\frac{u_{m}}{v_{m}}-1\right)t_{m}-\frac{w_{m}}{v_{m}}$ (111)
$=0,$ (112)
where (112) follows from (84). Thus, for all $m<\frac{n}{k}$,
$q_{m}-q_{m-1}$ $=\left(\frac{u_{m}}{v_{m}}-1\right)q_{m-1}-\frac{w_{m}}{v_{m}}\leq 0.$ (113)
Therefore, $f_{1}(m,q_{m-1})\leq q_{m-1}$. ∎
Figure 3: The three cases in the proof of Lemma 4.
Lemma 4.
For $\epsilon>0,k\geq 2,\text{ and }n>k$, for all $m$,
$f_{2}(q_{m-1})\leq q_{m-1}.$ (114)
Proof:
For some fixed $n$, $k$ and $\epsilon$, the function
$f_{2}(q)=\frac{v^{\prime}}{u^{\prime}}q+\frac{w^{\prime}}{u^{\prime}}$ (115)
is linear in $q$, with slope $\frac{v^{\prime}}{u^{\prime}}$ and intercept $\frac{w^{\prime}}{u^{\prime}}$. The slope
$\frac{v^{\prime}}{u^{\prime}}=\frac{e^{\epsilon}\left(\frac{1}{k}-1\right)}{-1+\frac{1}{k}-\frac{1}{n}}=\frac{e^{\epsilon}\left(\frac{1}{k}-1\right)}{\left(\frac{1}{k}-1\right)-\frac{1}{n}}$ (116)
is positive and less than $1$ for $0<\epsilon<\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. For $\epsilon>\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$, the slope is positive and greater than $1$. The intercept
$\frac{w^{\prime}}{u^{\prime}}=\frac{e^{\epsilon}-1-\frac{1}{n}}{-1+\frac{1}{k}-\frac{1}{n}}$ (117)
is positive for $\epsilon<\log(1+\frac{1}{n})$ and negative for $\epsilon>\log(1+\frac{1}{n})$. Thus, we can consider three regions of $\epsilon$ values as illustrated in Figure 3. Let $q^{*}$ denote the fixed point of $f_{2}$, i.e., the value of $q$ where $f_{2}(q)=q$. Thus
$q^{*}$ $=\frac{\frac{w^{\prime}}{u^{\prime}}}{1-\frac{v^{\prime}}{u^{\prime}}}$ (118)
$=\frac{w^{\prime}}{u^{\prime}-v^{\prime}}.$ (119)
Substituting (62)–(64) into (119) and simplifying, we have
$q^{*}$ $=\frac{e^{\epsilon}-1-\frac{1}{n}}{-1+\frac{1}{k}-\frac{1}{n}-e^{\epsilon}\left(\frac{1}{k}-1\right)}$ (120)
$=\frac{ne^{\epsilon}-n-1}{ne^{\epsilon}-n-1-\frac{n}{k}(e^{\epsilon}-1)}$ (121)
$=\frac{1}{1+\frac{\frac{n}{k}(e^{\epsilon}-1)}{1-n(e^{\epsilon}-1)}}.$ (122)
Recall that $q_{0}=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}$. We now consider each range of $\epsilon$ values and prove that $f_{2}(q_{m-1})\leq q_{m-1}$ holds for each case.
Case I: $0<\epsilon<\log(1+\frac{1}{n})$. In this region, we have $0<\frac{v^{\prime}}{u^{\prime}}<1$ and a positive intercept $\frac{w^{\prime}}{u^{\prime}}$. Thus, $q^{*}$ is positive. Moreover, $f_{2}(q)\leq q$ for $q\geq q^{*}$, so to prove the lemma it is enough to show that $q_{m}\geq q^{*}$ for all $m$. We will prove this by induction. For the base case, consider $m=0$. From (122), we observe that $0<1-n(e^{\epsilon}-1)<1$ in this range of $\epsilon$, which implies $q^{*}<q_{0}$. This proves the base case. Now, for the inductive step assume that $q_{m-1}\geq q^{*}$. Thus we have
$q_{m}=\max\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\}\geq f_{2}(q_{m-1})\geq q^{*}.$ (123)
This completes the proof for this case.
Case II: $\log(1+\frac{1}{n})\leq\epsilon\leq\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. In this region, the slope satisfies $0<\frac{v^{\prime}}{u^{\prime}}\leq 1$, but the intercept satisfies $\frac{w^{\prime}}{u^{\prime}}\leq 0$. Consequently, $f_{2}(q)\leq q$ for all $q\geq 0$, which proves the lemma in this case since by construction $q_{m-1}\geq 0$ for all $m$.
Case III: $\epsilon>\log\left(1+\frac{1}{n-\frac{n}{k}}\right)$. In this region, we have $\frac{v^{\prime}}{u^{\prime}}>1$ and a negative intercept $\frac{w^{\prime}}{u^{\prime}}$. Thus, $q^{*}$ is positive. Since $f_{2}(q)\leq q$ for $q\leq q^{*}$, it is enough to show that $q_{m}\leq q^{*}$ for all $m$. We prove this by induction. For the base case, consider $m=0$. We observe that $1-n(e^{\epsilon}-1)<0$ in this range of $\epsilon$. This implies that
$1+\frac{n}{k}(e^{\epsilon}-1)$ $>1+\frac{\frac{n}{k}(e^{\epsilon}-1)}{1-n(e^{\epsilon}-1)}.$ (124)
Therefore, $q_{0}<q^{*}$. Now, for the inductive step, assume for all $m^{\prime}<m$, $q_{m^{\prime}}\leq q^{*}$. Thus, $f_{2}(q_{m^{\prime}})\leq q_{m^{\prime}}$ for all $m^{\prime}<m$. From Lemma 3, we also know that $f_{1}(m^{\prime},q_{m^{\prime}-1})\leq q_{m^{\prime}-1}$ for all $m^{\prime}$. Thus, for all $m^{\prime}\leq m$,
$q_{m^{\prime}}=\max\{0,f_{1}(m^{\prime},q_{m^{\prime}-1}),f_{2}(q_{m^{\prime}-1})\}\leq\max\{0,q_{m^{\prime}-1}\}=q_{m^{\prime}-1}.$ (125)
That is, $q_{m^{\prime}}$ is non-increasing in $m^{\prime}$ for all $m^{\prime}\leq m$, which means $q_{m}\leq q_{0}$. As we have already shown that $q_{0}\leq q^{*}$, this proves that $q_{m}\leq q^{*}$, which completes the proof of the lemma. ∎
Lemma 5.
For $\epsilon>0,k\geq 2,n>k$, $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1\right\}$, and $u_{m},v_{m}$ and $w_{m}$ as defined in (65)–(67), the function $q_{m}$ satisfies the inequality condition
$u_{m}q_{m+1}<v_{m}q_{m}+w_{m}.$ (126)
Proof:
Using (65)–(67), we have, for $m=0$,
$u_{0}=\frac{1}{k}-\frac{1}{n},v_{0}=\frac{e^{\epsilon}}{k},w_{0}=-\frac{1}{n}.$ (127)
Then, we can write
$u_{0}q_{1}-v_{0}q_{0}$ $=u_{0}q_{1}-u_{0}q_{0}+u_{0}q_{0}-v_{0}q_{0}$
$=u_{0}(q_{1}-q_{0})+\left(\frac{1}{k}-\frac{1}{n}\right)q_{0}-\frac{e^{\epsilon}}{k}q_{0}$ (128)
$=u_{0}(q_{1}-q_{0})+\left(\frac{1}{k}+w_{0}\right)q_{0}-\frac{e^{\epsilon}}{k}q_{0}$ (129)
$=u_{0}(q_{1}-q_{0})+\frac{1}{k}q_{0}+w_{0}q_{0}-\frac{e^{\epsilon}}{k}q_{0}$ (130)
$=w_{0}q_{0}-\frac{e^{\epsilon}-1}{k}q_{0}-u_{0}(q_{0}-q_{1})$ (131)
$<w_{0}.$ (132)
Here, (132) follows from the fact that $0\leq q_{0}\leq 1$ and $q_{m}$ is non-increasing.
For $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1$, we can write
$u_{m}q_{m+1}$- $v_{m}q_{m}=u_{m}q_{m+1}-u_{m}q_{m-1}+w_{m}$ (133)
$=w_{m}+u_{m}(q_{m+1}-q_{m-1})$ (134)
$=w_{m}+\left(-\frac{m}{n}+\frac{1}{k}-\frac{1}{n}\right)(q_{m+1}-q_{m-1})$ (135)
$=w_{m}-\left(\frac{1}{k}-\frac{m+1}{n}\right)(q_{m-1}-q_{m+1})$ (136)
$<w_{m}.$ (137)
Here, (137) follows from the fact that $q_{m}$ is non-increasing and $\frac{1}{k}-\frac{m+1}{n}$ is non-negative for $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor-1$. Finally, for $q_{m+1}=0$, we have
$-v_{m}q_{m}$ $=w_{m}-\left(\frac{1}{k}-\frac{m+1}{n}\right)q_{m-1}$ (138)
$<w_{m}$ (139)
$\Rightarrow q_{m}$ $>-\frac{w_{m}}{v_{m}}.$ (140)
It follows from (108) that the right side of (140) is negative. Thus, (140) is a valid inequality, and we can conclude that $q_{m}$ satisfies (126) for the relevant range of $m$. ∎
Lemma 6.
For $\epsilon>0,k\geq 2,n>k$, $m=0,1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor$, and $u^{\prime},v^{\prime}$, and $w^{\prime}$ as defined in (62)–(64), the function $q_{m}$ satisfies the inequality condition
$(u^{\prime}-v^{\prime})q_{m}<w^{\prime}.$ (141)
Proof:
$(u^{\prime}-v^{\prime})q_{m}$ $=\left(-1+\frac{1}{k}-\frac{1}{n}-\frac{e^{\epsilon}}{k}+e^{\epsilon}\right)q_{m}$ (142)
$=\left(e^{\epsilon}-1-\frac{1}{n}+\frac{1}{k}(1-e^{\epsilon})\right)q_{m}$ (143)
$=w^{\prime}q_{m}+\frac{q_{m}}{k}(1-e^{\epsilon})$ (144)
$<w^{\prime}.$ (145)
Here, (139) follows from the fact that $0\leq q_{m}\leq 1$ and $\frac{q_{m}}{k}(1-e^{\epsilon})<0$. ∎
Lemma 7.
For $\epsilon>0,k\geq 2,n>k$, $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor$, and $u^{\prime},v^{\prime}$, and $w^{\prime}$ as defined in (62)–(64), the function $q_{m}$ satisfies the inequality condition
$u^{\prime}q_{m-1}<v^{\prime}q_{m}+w^{\prime}.$ (146)
Proof:
$u^{\prime}q_{m-1}-v^{\prime}q_{m}$ $=u^{\prime}q_{m-1}-v^{\prime}q_{m-1}+v^{\prime}q_{m-1}-v^{\prime}q_{m}$ (147)
$=(u^{\prime}-v^{\prime})q_{m-1}+v^{\prime}(q_{m-1}-q_{m})$ (148)
$=(u^{\prime}-v^{\prime})q_{m-1}+e^{\epsilon}\left(\frac{1}{k}-1\right)(q_{m-1}-q_{m})$ (149)
$<(u^{\prime}-v^{\prime})q_{m-1}$ (150)
$<w^{\prime}.$ (151)
Here, (144) follows from the fact that $e^{\epsilon}\left(\frac{1}{k}-1\right)<0$ and $q_{m}$ is non-increasing, and (145) follows from Lemma 6. Therefore,
$u^{\prime}q_{m-1}<v^{\prime}q_{m}+w^{\prime}.$ (152)
Appendix C Proof of Theorem 3
The utility of DS-ROO is measured in terms of the total variation distance between the output distribution $Q$ of Algorithm 2 and the unknown true distribution $P$. Hence, to prove Theorem 3, we derive an upper bound on $d_{TV}(P,Q)$. The proof is structured around two auxiliary lemmas. Recall that we define the minimum number of occurrences of an element of the alphabet as $M=n\cdot\min_{x\in\mathcal{X}}\hat{P}_{X^{n}}(x)$. In Lemma 8, we show that under the assumption $P(x)\geq\gamma$ for all $x\in[k]$, the probability of $\min_{x\in\mathcal{X}}\hat{P}_{X^{n}}(x)$ falling below a threshold $\gamma^{\prime}<\gamma$ decays exponentially in dataset size $n$. We then decompose QPQ-P by conditioning on the two events $M\geq m_{0}$ and $M<m_{0}$, where $m_{0}=n\gamma^{\prime}$, and bound the contribution of each case separately. In Lemma 9, we derive conditions under which the data-specific obscuring probability $q_{m_{0}}=0$. Finally, we combine these results to obtain the upper bound on $d_{TV}(P,Q)$, which completes the proof.
Now, using Lemma 8, we can write
$\Pr\{M<m_{0}\}$ $\leq ke^{-2n(\gamma-\gamma^{\prime})^{2}},$ (153)
where $m_{0}=n\gamma^{\prime}$, and $\gamma^{\prime}<\gamma$. Let $\pi=\Pr\{M<m_{0}\}$ and $\delta=ke^{-2n(\gamma-\gamma^{\prime})^{2}}$. Recall that the output distribution is
$Q(y)=\mathbb{E}_{X^{n}}\left[\frac{q_{M}}{k}+(1-q_{M})\hat{P}_{X^{n}}(y)\right].$ (154)
We have
$Q(y)-P(y)$ $=\mathbb{E}_{X^{n}}\left[\frac{q_{M}}{k}+(1-q_{M})\hat{P}_{X^{n}}(y)-\hat{P}_{X^{n}}(y)\right]$ (155)
$=\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\right]$ (156)
$=\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right](1-\pi)$
$\hskip 56.9055pt+\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\pi.$ (157)
Now, we can write
$\sum_{y\in\mathcal{X}}\left|Q(y)-P(y)\right|$
$\quad=\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right](1-\pi)$
$\quad\quad+\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\pi\Bigg|$ (158)
$\quad\leq(1-\pi)\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right]\Bigg|$
$\quad\quad+\pi\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\Bigg|$ (159)
$\quad\leq\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right]\Bigg|$
$\quad\quad+\delta\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\Bigg|.$ (160)
Let
$A$ $=\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right]\Bigg|,$ (161)
$B$ $=\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\Bigg|.$ (162)
We can bound $A$ as follows:
$A$ $=\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M\geq m_{0}\right]\Bigg|$ (163)
$\leq\sum_{y\in\mathcal{X}}\mathbb{E}_{X^{n}}\left[q_{M}\left|\frac{1}{k}-\hat{P}_{X^{n}}(y)\right|\Bigg|M\geq m_{0}\right]$ (164)
$=\mathbb{E}_{X^{n}}\left[q_{M}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-\hat{P}_{X^{n}}(y)\right|\Bigg|M\geq m_{0}\right]$ (165)
$\leq q_{m_{0}}\mathbb{E}_{X^{n}}\left[\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-\hat{P}_{X^{n}}(y)\right|\Bigg|M\geq m_{0}\right]$ (166)
$=2q_{m_{0}}\mathbb{E}_{X^{n}}\left[d_{TV}\left(\frac{1}{k},\hat{P}_{X^{n}}(y)\right)\Bigg|M\geq m_{0}\right]$ (167)
$\leq 2q_{m_{0}}\left(1-\frac{1}{k}\right).$ (168)
Here, (166) follows from the fact that for $M\geq m_{0}$, $q_{M}\leq q_{m_{0}}$ since $q_{M}$ is non-decreasing in $M$ as proved in Lemmas 3 and 4. Thus, we have
$\mathbb{E}_{X^{n}}[q_{M}|M\geq m_{0}]\leq q_{m_{0}}.$ (169)
Lastly, (168) follows from the upper bound on $d_{TV}\left(\frac{1}{k},\hat{P}_{X^{n}}(y)\right)$ as illustrated in the proof of Theorem 1. We can bound $B$ in a similar manner:
$B$ $=\sum_{y\in\mathcal{X}}\Bigg|\mathbb{E}_{X^{n}}\left[q_{M}\left(\frac{1}{k}-\hat{P}_{X^{n}}(y)\right)\bigg|M<m_{0}\right]\Bigg|$ (170)
$\leq\sum_{y\in\mathcal{X}}\mathbb{E}_{X^{n}}\left[q_{M}\left|\frac{1}{k}-\hat{P}_{X^{n}}(y)\right|\Bigg|M<m_{0}\right]$ (171)
$=\mathbb{E}_{X^{n}}\left[q_{M}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-\hat{P}_{X^{n}}(y)\right|\Bigg|M<m_{0}\right]$ (172)
$=2\mathbb{E}_{X^{n}}\left[q_{M}\cdot d_{TV}\left(\frac{1}{k},\hat{P}_{X^{n}}(y)\right)\Bigg|M<m_{0}\right]$ (173)
$\leq 2q_{0}\left(1-\frac{1}{k}\right).$ (174)
Now, combining these, we have
$\sum_{y\in\mathcal{X}}\left|Q(y)-P(y)\right|\leq A+\delta B\leq 2(q_{m_{0}}+q_{0}\delta)\left(1-\frac{1}{k}\right).$ (175)
The total variation distance between the input distribution $P$ and output distribution $Q$ is thus
$d_{TV}(P,Q)$ $=\frac{1}{2}\sum_{y\in\mathcal{X}}\left|Q(y)-P(y)\right|$ (176)
$\leq\left(q_{m_{0}}+q_{0}ke^{-2n(\gamma-\gamma^{\prime})^{2}}\right)\left(1-\frac{1}{k}\right).$ (177)
We achieve the tightest upper bound on $d_{TV}$ when $q_{m_{0}}=0$, i.e., when Algorithm 2 never obscures the empirical distribution. Lemma 9 shows that $q_{m_{0}}=0$ if
$m_{0}\geq\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\}.$ (178)
The total variation distance is then
$d_{TV}(P,Q)\leq\alpha$ $=q_{0}ke^{-2n(\gamma-\gamma^{\prime})^{2}}\left(1-\frac{1}{k}\right)$ (179)
$=q_{0}(k-1)e^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}.$ (180)
C-A Auxiliary Lemmas
Lemma 8.
If $P(x)\geq\gamma$ for all $x\in\mathcal{X}$ where $\mathcal{X}=\{1,2,\dots,k\}$, then for some $\gamma^{\prime}<\gamma$,
$\Pr\left\{\min_{x\in\mathcal{X}}\hat{P}_{X^{n}}(x)<\gamma^{\prime}\right\}\leq ke^{-2n(\gamma-\gamma^{\prime})^{2}}.$ (181)
Proof:
$\Pr\left\{\min_{x\in\mathcal{X}}\hat{P}_{X^{n}}(x)<\gamma^{\prime}\right\}$
$\leq\sum_{x\in\mathcal{X}}\Pr\left\{\hat{P}_{X^{n}}(x)<\gamma^{\prime}\right\}$ (182)
$=\sum_{x\in\mathcal{X}}\Pr\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)<\gamma^{\prime}\right\}$ (183)
$=\sum_{x\in\mathcal{X}}\Pr\left\{\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)<n\gamma^{\prime}\right\}$ (184)
$=\sum_{x\in\mathcal{X}}\Pr\Bigg\{\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)-\mathbb{E}\left[\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)\right]$
$\quad\hskip 40.0pt<n\gamma^{\prime}-\mathbb{E}\left[\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)\right]\Bigg\}$ (185)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{x\in\mathcal{X}}\Pr\Bigg\{\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)-\mathbb{E}\left[\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)\right]$
$\quad\hskip 40.0pt<n\gamma^{\prime}-n\gamma\Bigg\}$ (186)
$=\sum_{x\in\mathcal{X}}\Pr\Bigg\{\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)-\mathbb{E}\left[\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)\right]$
$\quad\hskip 40.0pt<-n(\gamma-\gamma^{\prime})\Bigg\}$ (187)
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{x\in\mathcal{X}}e^{-2n(\gamma-\gamma^{\prime})^{2}}$ (188)
$=ke^{-2n(\gamma-\gamma^{\prime})^{2}}$ (189)
Here, $(a)$ follows from our assumption that $P(x)=\Pr\{X_{i}=x\}\geq\gamma$ for all $x$. Then, $\mathbb{E}\left[\sum_{i=1}^{n}\mathbf{1}(X_{i}=x)\right]=\sum_{i=1}^{n}\mathbb{E}\left[\mathbf{1}(X_{i}=x)\right]=\sum_{i=1}^{n}\Pr\{X_{i}=x\}\geq n\gamma$. Finally, we apply Hoeffding’s inequality in $(b)$. Note that, since $(\gamma-\gamma^{\prime})^{2}$ is always positive, the right side of the inequality goes to 0 as the sample size $n$ goes to $\infty$. ∎
Lemma 9.
Given $k\geq 2$, $\epsilon>0$, and
$q_{0}=\frac{1}{1+\frac{n}{k}\left(e^{\epsilon}-1\right)},$ (190)
if there exists an integer $m_{0}$ such that
$m_{0}$ $\geq\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\},$ (191)
$n$ $\geq\max\left\{\frac{m_{0}}{\gamma},\frac{k-1+\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}}{2(e^{\epsilon}-1)^{2}},2k\right\},$ (192)
then the corresponding data-specific obscuring probability is zero, i.e.,
$q_{m_{0}}=0.$ (193)
Proof:
For some $m_{0}$, we can find the corresponding $q_{m_{0}}$ using the following formula stated in Algorithm 2,
$q_{m_{0}}$ $=\max\left\{0,\frac{u_{m_{0}}}{v_{m_{0}}}q_{{m_{0}}-1}-\frac{w_{m_{0}}}{v_{m_{0}}},\frac{v^{\prime}}{u^{\prime}}q_{{m_{0}}-1}+\frac{w^{\prime}}{u^{\prime}}\right\}.$ (194)
We can simplify the second recursive expression in (194) as follows:
$\frac{v^{\prime}}{u^{\prime}}q_{0}+\frac{w^{\prime}}{u^{\prime}}$ $=\frac{e^{\epsilon}(\frac{1}{k}-1)}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}+\frac{e^{\epsilon}-1-\frac{1}{n}}{-1+\frac{1}{k}-\frac{1}{n}}$ (195)
$=\frac{e^{\epsilon}(\frac{1}{k}-1)}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}+\frac{(e^{\epsilon}-1-\frac{1}{n})(1+\frac{n}{k}(e^{\epsilon}-1))}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}$ (196)
$=\frac{e^{\epsilon}(\frac{1}{k}-1)+(e^{\epsilon}-1-\frac{1}{n})(1+\frac{n}{k}(e^{\epsilon}-1))}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}$ (197)
$=\frac{\frac{e^{\epsilon}}{k}-e^{\epsilon}+e^{\epsilon}+e^{\epsilon}\frac{n}{k}(e^{\epsilon}-1)-1-\frac{n}{k}(e^{\epsilon}-1)-\frac{1}{n}-\frac{e^{\epsilon}}{k}+\frac{1}{k}}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}$ (198)
$=\frac{-1+\frac{1}{k}-\frac{1}{n}+\frac{n}{k}(e^{\epsilon}-1)^{2}}{-1+\frac{1}{k}-\frac{1}{n}}q_{0}$ (199)
$=\left(1+\frac{\frac{n}{k}(e^{\epsilon}-1)^{2}}{-1+\frac{1}{k}-\frac{1}{n}}\right)q_{0}$ (200)
$=\left(1+\frac{n^{2}(e^{\epsilon}-1)^{2}}{-n(k-1)-k}\right)q_{0}.$ (201)
If the right side of (201) is negative, the second recursion in (194) is never chosen as $q_{m_{0}}$. This happens when
$1+\frac{n^{2}(e^{\epsilon}-1)^{2}}{-n(k-1)-k}\leq 0.$ (202)
Observe that the denominator in (202) is negative. Then we can write the inequality as
$n^{2}(e^{\epsilon}-1)^{2}-n(k-1)-k\geq 0.$ (203)
Solving for $n$, we have
$n\geq\frac{k-1\pm\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}}{2(e^{\epsilon}-1)^{2}}.$ (204)
Since $k-1<\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}$, the valid bound on the dataset size is
$n\geq\frac{k-1+\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}}{2(e^{\epsilon}-1)^{2}}.$ (205)
When the dataset size satisfies the above lower bound, the only term to consider from (194) is the first recursive expression. We have
$\frac{u_{m_{0}}}{v_{m_{0}}}q_{{m_{0}}-1}-\frac{w_{m_{0}}}{v_{m_{0}}}$ $=q_{0}\prod_{i=1}^{m_{0}}\frac{u_{i}}{v_{i}}-\sum_{j=1}^{m_{0}-1}\left(\frac{w_{j}}{v_{j}}\prod_{i=j+1}^{m_{0}}\frac{u_{i}}{v_{i}}\right)-\frac{w_{m_{0}}}{v_{m_{0}}}=q_{0}\prod_{i=1}^{m_{0}}\frac{u_{i}}{v_{i}}+\mathcal{I}.$ (206)
We simplify the coefficient of $q_{0}$ as follows:
$\prod_{i=1}^{m_{0}}\frac{u_{i}}{v_{i}}$ $=\prod_{i=1}^{m_{0}}\frac{\frac{1}{k}-\frac{i}{n}-\frac{1}{n}}{e^{\epsilon}\left(\frac{1}{k}-\frac{i}{n}\right)}=\prod_{i=1}^{m_{0}}e^{-\epsilon}\frac{\frac{1}{k}-\frac{i+1}{n}}{\frac{1}{k}-\frac{i}{n}}=e^{-m_{0}\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{1}{n}}<e^{-m_{0}\epsilon}.$ (207)
To analyze the remaining terms, we can write
$\mathcal{I}$ $=-\sum_{j=1}^{m_{0}-1}\left(\frac{w_{j}}{v_{j}}\prod_{i=j+1}^{m_{0}}\frac{u_{i}}{v_{i}}\right)-\frac{w_{m_{0}}}{v_{m_{0}}}$ (208)
$=-\sum_{j=1}^{m_{0}-1}\left(\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}\right)-\frac{w_{m_{0}}}{v_{m_{0}}}$ (209)
$=-\sum_{j=1}^{m_{0}}\frac{w_{j}}{v_{j}}\underbrace{e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}}_{\text{positive and less than 1}}.$ (210)
The terms $\frac{w_{j}}{v_{j}}$ has the general form
$\frac{w_{j}}{v_{j}}$ $=\frac{\frac{j}{n}(e^{\epsilon}-1)-\frac{1}{n}}{e^{\epsilon}\left(\frac{1}{k}-\frac{j}{n}\right)},$ (211)
which is increasing in $j$, with negative values for $j=1,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor$ and positive values for $j=\left\lceil\frac{1}{e^{\epsilon}-1}\right\rceil,\dots,m_{0}$. Therefore, we can split the summation into two terms
$\mathcal{I}$ $=-\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}-\sum_{j=\left\lceil\frac{1}{e^{\epsilon}-1}\right\rceil}^{m_{0}}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}$ (212)
$=\mathcal{I_{+}}+\mathcal{I_{-}},$ (213)
where
$\mathcal{I_{+}}$ $=-\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}},$ (214)
$\mathcal{I_{-}}$ $=-\sum_{j=\left\lceil\frac{1}{e^{\epsilon}-1}\right\rceil}^{m_{0}}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}.$ (215)
Note that the terms in $\mathcal{I_{-}}$ appear in (212) only if $m_{0}$ is at least $\left\lceil\frac{1}{e^{\epsilon}-1}\right\rceil$. Hence, we derive an upper bound on $\mathcal{I}$ under the assumption $m_{0}>\frac{2}{e^{\epsilon}-1}$. To this end, we first find an upper bound on $\mathcal{I_{+}}$ and a lower bound on $|\mathcal{I_{-}}|$. We have
$\mathcal{I_{+}}$ $=-\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}$ (216)
$\leq\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}-\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}$ (217)
$\leq\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}-\frac{w_{1}}{v_{1}}e^{-(m_{0}-j)\epsilon}$ (218)
$=-\frac{w_{1}}{v_{1}}e^{-m_{0}\epsilon}\sum_{j=1}^{\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}e^{j\epsilon}$ (219)
$=\frac{-\frac{1}{n}\left(e^{\epsilon}-1\right)+\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{1}{n})}e^{-m_{0}\epsilon}e^{\epsilon}\frac{e^{\epsilon\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}-1}{e^{\epsilon}-1}$ (220)
$=\frac{k(2-e^{\epsilon})}{n-k}e^{-m_{0}\epsilon}\frac{e^{\epsilon\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor}-1}{e^{\epsilon}-1}$ (221)
$\leq\frac{k(2-e^{\epsilon})}{n-k}e^{-m_{0}\epsilon}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}.$ (222)
Here we used the fact that each $\frac{w_{j}}{v_{j}}$ term in the sum in (217) is upper bounded by $\frac{w_{1}}{v_{1}}$. We can lower bound $|\mathcal{I_{-}}|$ as follows:
$|\mathcal{I_{-}}|$ $=\sum_{j=\left\lceil\frac{1}{e^{\epsilon}-1}\right\rceil}^{m_{0}}\frac{w_{j}}{v_{j}}e^{-(m_{0}-j)\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{j+1}{n}}$ (223)
$\geq\frac{w_{m_{0}}}{v_{m_{0}}}e^{-(m_{0}-m_{0})\epsilon}\frac{\frac{1}{k}-\frac{m_{0}+1}{n}}{\frac{1}{k}-\frac{m_{0}+1}{n}}$ (224)
$=\frac{w_{m_{0}}}{v_{m_{0}}}=\frac{\frac{m_{0}}{n}\left(e^{\epsilon}-1\right)-\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{m_{0}}{n})}.$ (225)
Combining (222) and (225), we have
$\mathcal{I}$ $\leq e^{-m_{0}\epsilon}\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{\frac{m_{0}}{n}\left(e^{\epsilon}-1\right)-\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{m_{0}}{n})}$ (226)
$\leq e^{-m_{0}\epsilon}\left(\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{\frac{m_{0}}{n}\left(e^{\epsilon}-1\right)-\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{m_{0}}{n})}e^{m_{0}\epsilon}\right).$ (227)
Thus, we have the upper bound on (206),
$\frac{u_{m_{0}}}{v_{m_{0}}}q_{{m_{0}}-1}-\frac{w_{m_{0}}}{v_{m_{0}}}$ $=q_{0}\prod_{i=1}^{m_{0}}\frac{u_{i}}{v_{i}}+\mathcal{I}\leq e^{-m_{0}\epsilon}(q_{0}+C),$ (228)
where
$C$ $=\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{\frac{m_{0}}{n}\left(e^{\epsilon}-1\right)-\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{m_{0}}{n})}e^{m_{0}\epsilon}.$ (229)
Since $q_{m_{0}}$ is monotonic and non-increasing, (228) is a universal upper bound on $q_{m_{0}}$. Now, $q_{m_{0}}$ will be zero if the right side of (228) is non-positive, i.e.,
$e^{-m_{0}\epsilon}(q_{0}+C)\leq 0.$ (230)
Since $e^{-m_{0}\epsilon}>0$ for the relevant range of $m_{0}$ and $\epsilon$, we require
$q_{0}+C\leq 0.$ (231)
Substituting the expressions for $q_{0}$ and $C$ into the left side of (231), we have
$q_{0}+C$ $=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}+\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{\frac{m_{0}}{n}\left(e^{\epsilon}-1\right)-\frac{1}{n}}{e^{\epsilon}(\frac{1}{k}-\frac{m_{0}}{n})}e^{m_{0}\epsilon}$ (232)
$\leq\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}+\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{\frac{1}{n}}{e^{\epsilon}\left(\frac{1}{k}-\frac{1}{n}\right)}e^{m_{0}\epsilon}$ (233)
$=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}+\frac{k(2-e^{\epsilon})}{n-k}\frac{e^{\epsilon\frac{1}{e^{\epsilon}-1}}-1}{e^{\epsilon}-1}-\frac{k}{n-k}e^{(m_{0}-1)\epsilon}$ (234)
$\leq\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}+\frac{k}{n-k}\frac{(2-1-\epsilon)\left(e^{\frac{\epsilon}{1+\epsilon-1}}-1\right)}{1+\epsilon-1}-\frac{k}{n-k}e^{(m_{0}-1)\epsilon}$ (235)
$=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}+\frac{k}{n-k}\frac{(1-\epsilon)(e-1)}{\epsilon}-\frac{k}{n-k}e^{(m_{0}-1)\epsilon}$ (236)
$\leq\frac{1}{1+\frac{n}{k}\epsilon}+\frac{k}{n-k}\frac{(1-\epsilon)(e-1)}{\epsilon}-\frac{k}{n-k}e^{(m_{0}-1)\epsilon}$ (237)
$=\frac{k}{n\epsilon+k}+\frac{k}{n-k}\frac{(1-\epsilon)(e-1)}{\epsilon}-\frac{k}{n-k}e^{(m_{0}-1)\epsilon}$ (238)
$=\frac{k}{n-k}\left(\frac{n-k}{n\epsilon+k}+\frac{(1-\epsilon)(e-1)}{\epsilon}-e^{(m_{0}-1)\epsilon}\right)$ (239)
$=\frac{k}{n-k}\left(\frac{1}{\epsilon}\frac{n\epsilon-k\epsilon}{n\epsilon+k}+\frac{(1-\epsilon)(e-1)}{\epsilon}-e^{(m_{0}-1)\epsilon}\right)$ (240)
$\leq\frac{k}{n-k}\left(\frac{1}{\epsilon}+\frac{(1-\epsilon)(e-1)}{\epsilon}-e^{(m_{0}-1)\epsilon}\right)$ (241)
$<\frac{1}{\epsilon}+\frac{(1-\epsilon)(e-1)}{\epsilon}-e^{(m_{0}-1)\epsilon}$ (242)
$=\frac{1}{\epsilon}+\frac{e-e\epsilon-1+\epsilon}{\epsilon}-e^{(m_{0}-1)\epsilon}$ (243)
$=\frac{1}{\epsilon}+\frac{e}{\epsilon}-e-\frac{1}{\epsilon}+1-e^{(m_{0}-1)\epsilon}$ (244)
$=\frac{e}{\epsilon}-e+1-e^{(m_{0}-1)\epsilon}.$ (245)
Here, (233) follows from the fact that we assume $m_{0}>\frac{2}{e^{\epsilon}-1}$ to obtain the tightest upper bound on $q_{m_{0}}$. In (235) and (237), we use the bound $e^{\epsilon}\geq 1+\epsilon$ from the Taylor expansion of $e^{\epsilon}$. Finally, (241) and (242) follow from the assumption that $n>2k$. Thus, the condition in (231) holds if
$\frac{e}{\epsilon}-e+1-e^{(m_{0}-1)\epsilon}\leq 0.$ (246)
Rearranging (246), we obtain
$m_{0}\geq 1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right).$ (247)
In the low-privacy regime (larger $\epsilon$), this bound requires a smaller $m_{0}$, while stronger privacy demands larger $m_{0}$. ∎
Acknowledgment
This work is supported in part by NSF grants CIF-1901243, CIF-2312666, and CIF-2007688.